Solve (x + 2)y"-(2x + 5)y'

+ 2y = (x + 1)ex


1
Expert's answer
2021-07-29T16:18:12-0400

Solution

One of the solutions of homogeneous equation (x + 2)y0"-(2x + 5)y0'+ 2y0 = 0 let’s find in the form y0(x) = erx. Substitution into equation gives

(x + 2)r2-(2x + 5)r+ 2= 0   =>   x(r2-2r)+(2r2-5r+2) = 0  => from first brackets equal to zero r=0, r=2 and from second r = 1/2, r = 2  =>  r = 2

So y01(x) = e2x.

Once we have this first solution we will then assume that a second solution of homogeneous equation will have the form

y(x) = y01(x)u(x)

Plugging this into the homogeneous equation gives

(x+2)u’’+(2x+3)u’ = 0

New function w(x) = u’(x)  => (x+2)w’ = -(2x+3)w  =>  dw/w = -(2x+3)dx/(x+2)  => ln|w| = - 2x+ln|x+2| + C1  =>  w = C1 (x+2)e-2x  => u(x) = C2+∫w(x)dx = C2+ C1∫(x+2)e-2x dx = C2 - C1 (x/2+5/4) e-2x    =>   y0(x) = y01(x)u(x) = C2 e2x - C1 (x/2+5/4)    

Partial solution of initial nonhomogeneous equation let’s find in the form y1(x) = Aex  

Plugging this into the nonhomogeneous equation gives

A[(x+2)-(2x-5)+2] = x+1  => A(-x-1) = x+1  = >   A = -1  =>  y1(x) = - ex    

Therefore solution of initial equation is

y(x) = y0(x) + y1(x) = C2 e2x - C1 (x/2+5/4) - ex